answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

A department store receives a shipment of 23 new portable radios There are 5 defective radios in the shipment If 4 radios are selected for display What is the probability that 2 of them are defective?

It is possible to choose 4 radios out of 23 in (23 choose 4)=8855 ways. The number of those combinations where there are exactly 2 indefective and 2 defective is: (18 (the number of those indefective) choose 2)*(5 (the number of defective radios) choose 2)=1530 divided by 8855=1530/8855=306/1771 which is 17.2784%.

A group of 10 people consists of 5 men and 5 women A committee of 4 is chosen from this group What is the probability that one or more of the committee members is a man?

The probability that there are 1,2,3 or 4 men is 1-(the probability that no men are selected). First we select the first person. The probability that this person is a woman is 5/10=1/2. For second person it is 4/9, then 3/8 and finally 2/7. We multiply these together: (1*4*3*2)/(2*9*8*7)=24/1008. This is the probability that every single person in the committee is a woman. One minus that probability is 984/1008=41/42 which is 97.619%

Read more >> Options >>

http://www.answers.com?initiator=FFANS

What is probability how you use or solve probability?

Probability is the chance (in percentage or decimal) of a particular event to happen.

lets say that you tossed a coin. the possible events to happen are ending up with heads or tails. the probability of having a head is 50% or .5 while the probability of having a tails is 50% or .5.

to solve for the probability, divide the particular event with the total number of possible events.

ex. what is the probability of getting a 3 when you rolled a dice?

particular event= having a 3= 1 event

total number of events= having either a 1, 2, 3, 4, 5, or 6= 6 events

particular event/ total number of events= 1/6

hoped i helped!

What is a random variable in probability theory?

Random Variable in probability theory is defined as follows: Assuming you have variables Xi where i is an integer ie: i=1,2,3.......n a variable Xi is called a random variable iff(if and only iff) and random selection yields a variable Xi for i=1,2.........,n with the same likelihood of appearance. i.e prob(X=Xi)=1/n

What is the probability that 4 randomly selected people all have different birthdays?

Let us assume that there are exactly 365 days in a year and that birthdays are uniformly randomly distributed across those days.

First, what is the probability that 2 randomly selected people have different birthdays? The second person's birthday can be any day except the first person's, so the probability is 364/365.

What is the probability that 3 people will all have different birthdays? We already know that there is a 364/365 chance that the first two will have different birthdays. The third person must have a birthday that is different from the first two: the probability of this happening is 363/365. We need to multiply the probabilities since the events are independent; the answer for 3 people is thus 364/365 × 363/365.

You should now be able to solve it for 4 people.

What is the chances of winning the lottery?

It depends on the particular lottery. For Mega Millions, which is five balls out of a field of 56, and one ball out of a field of 46, the odds of hitting the jackpot are 1 in 56 * 55 * 54 * 53 * 52 / 5 / 4 / 3 / 2 * 46 or 1 in 175,711,536. In short - slim to none.

The odds are far greater for getting in a fatal automobile accident or being struck by lightning. Gee - I should play because I have been struck by lightning!

What is the probability of giving birth to a baby girl?

You have a 1 in 2 chance of having a girl. No matter how many boys you have had in the past or how many girls you have had in the past, each pregnancy presents the opportunity to have a girl or a boy with a 50% chance. The sperm does not know that it had already produced ,say, 5 boys and needs to produce a girl for you and your partner. So with every pregnancy there is a 50% chance you will have a girl and a 50% chance you will have a boy.

What is the probability of a 28-year-old female bearing a baby with Down syndrome?

If you look at the link below, the probability of a 28 year old female having a baby with Down syndrome is a little less than 1 in 1000.

What are the chances of rolling the number seven with two dice?

There are 36 permutations of two dice. Of these, six have a sum of seven. The probability, then, of rolling a seven on two dice is 6 in 36, or 1 in 6, or about 0.1667.

It does not matter if two dice are rolled one time, or if one die is rolled two times. The probability of rolling a sum of 7 is still about 0.1667.
The first thing to do is to identify all possible combinations which could give the sum of 7. These are: 6 and 1, 5 and 2, 4 and 3, 3 and 4, 2 and 5, and 1 and 6.

Thus there are 6 possible combinations.

For each of these combinations, the probability of getting the 1st number correct is 1/6 and the probability of getting the second number correct is 1/6, and so the overall probability of the combination is 1/36

There are 6 ways to get 7 though so 6x 1/36 = 1/6

Which means the probability that 2 dice rolls will add up to 7 is 1/6

What is the probability of getting three tails if you flip a coin three times?

Since the probability of getting tails is 50% or 0.5, the probability of three tails would be 0.5*0.5*0.5=0.125 or 12.5 %

What is the probability tossing a fair coin n times?

The probability will b 0.5. since a coin tossed n times has 2n no. of desired results, among which only half that is n times it can b tail. Hence P = n/2n i.e 0.5.

What three digit number has the most factors?

Using prime factorization, there are 2 three digit numbers with 9 factors:

512 and 768.

There are no three digit numbers with 10 or more prime factors.

512=2^9 and 768=3*2^8

The probability that an individual is left handed is 0.11 In a class of 40 students what is the probability of finding five left handers?

If the probability of a person being left handed is .11, then the probability of a person being right handed is .89 (1.00 - .11). In a class of 40 people, there are 40! (40 factorial) = 40 x 39 x 38 x .... x 3 x 2 x 1 possible Permutations of students. Within this number, there are many that will yield exactly 5 left handed people. 40-P-5 is the number of Permutations of 5 within a class of 40 (sorry, can't use the correct formating here!): = 40! / (40 - 5) ! = 40! / 35! = 40 x 39 x 38 x 37 x 36 Since we don't care about order..... we need to factor out the number of identical permutations. 40-C-5 = 40-P-5 / 5! So.... (40 x 39 x 38 x 37 x 36) / (5 x 4 x 3 x 2 x 1) = 658,008 Now, what is the probability of any one such combination? (.89 to the power 35) x (.11 to the power 5) = 0.000000273 So the probability of exactly 5 left handed people in the class is 658,008 x .000000273 = 17.9408137 %

The probability that Luis will pass his statistics test is 0.33 Find the probability that he will fail his statistics test?

Since Luis can only pass or fail his test, and we know that the probability of all possible outcomes is 1, we therefore see that P(Pass) + P(Fail) = 1

*where P(X) = the probability of event X*

So if P(Pass) = .33;

then P(Fail) = 1 - P(Pass)

= 1 - .33

= .66 (or 66% chance that Luis will fail)

What is the probability of getting head lice and how do you stop yourself from being scared of getting head lice?

If you have primary school age children then the likelihood of them getting head lice (repeatedly!( these days is very high - believe me! Chances are you will have them too at some point. Other than that, I would think your chances of catching them are quite low.

To stop yourself being scared the best thing to do is to realise that it happens to lots of normal people, and to lots of normal people's children. You get a more realistic picture when you realise that - it is not a major catastrophe, nor something that can't be fixed. They can even be eradicated naturally - it just takes patient combing with a fine tooth comb - for a few weeks to ensure none are left.

One way to keep them away, there is home remedy.. it conset of rosemary poppuri a lil of olive oil, (because you don't want to dry out the sclap with the poppuri) so u mix it all up! apply around the edges of the hair. where lice tend to enter.. for sum reason the lice don't like it... so there u have it good luck.. if this works for u let me no!! muah...

Why be scared lol it happens you just need to get the right stuff to put on your hair.... and...... bobs your uncle!!!!! well he's my dad but you get my point

How many 9 digit combinations are there for the numbers 1 through 9?

9 times 9 times 9 times 9 times 9 times 9 times 9 times 9 times 9 =

Imagine it's a combination lock with 9 tumblers. Each tumbler has 9 "positions" - the digits from 1 to 9. So you multiply the number of digits [9] against itself [9] for as many tumblers are in the combination (9).

Tumbler # 1 2 3 4 5 6 7 8 9

Number of

possible

digits 9 9 9 9 9 9 9 9 9

How does probability affect your life?

lol answer it for urself dont rely on this ,this is troll